40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16950     Entry time: Mon Jun 27 13:25:50 2022
Author: Cici 
Type: Update 
Category: General 
Subject: Characterizing the Transfer Loop 

[Deeksha, Cici]

We first took data of a simple low pass filter, and attempted to perform a fit to both the magnitude and phase in order to find the Z of the components. Once we felt confident in our ability to measure tranfer functions, we took data and plotted the transfer function of the existing control loop of the AUX laser. What we found generally followed the trend of, but was lower than, 10^4/f, which is what we hoped to match, and also had a strange unexplained notch ~1.3 kHz. The magnitude and phase data both got worse after around 40-50 kHz, which we believe is because the laser came out of lock near the end of the run. 

Edit: 

[Attachment 2 and 3] are the frequency response of the low pass filter, curves fitted using least squares in python.

[Attachment 1 and 4] is the same measurement of OLTF of the actual AUX circuit, and the control diagram pointing out the location of excitation and test point.

Attachment 1: TF_measurement_b.png  141 kB  Uploaded Mon Jun 27 14:34:05 2022  | Hide | Hide all
TF_measurement_b.png
Attachment 2: transfer_function_mag_fit.png  47 kB  Uploaded Mon Jun 27 14:36:45 2022  | Hide | Hide all
transfer_function_mag_fit.png
Attachment 3: transfer_function_phase_fit.png  57 kB  Uploaded Mon Jun 27 14:36:55 2022  | Hide | Hide all
transfer_function_phase_fit.png
Attachment 4: control_flow.png  68 kB  Uploaded Wed Jun 29 11:32:41 2022  | Hide | Hide all
control_flow.png
ELOG V3.1.3-